Circular No. 7882

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002cj IN ESO 582-G5 M. Ganeshalingam and W. D. Li, University of California at Berkeley,

report the LOTOSS (cf. IAUC 7514) discovery of an apparent supernova on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Apr. 21.5 (mag ∼ 16.0) and 22.5 UT (mag ∼ 15.9). SN 2002cj is located at α = 15^{h}21^{m}20^{s}.71, δ = −19^{o}51′27′′.3 (equinox 2000.0), which is 1′′.4 west and 3′′.9 south of the nucleus of ESO 582-G5. KAIT images taken on Apr. 9.5 (limiting mag ∼ 16.0) and Mar. 28.5 (limiting mag ∼ 18.0) showed nothing at this position.

V4741 SAGITTARII = POSSIBLE NOVA IN SAGITTARIUS N. N. Samus, Institute of Astronomy, Moscow; O. Durlevich, Sternberg

Institute; and E. Kazarovets, Institute of Astronomy, inform us that the designation V4741 Sgr has been given to this object (cf. IAUC 7878, 7879).

M. Fujii, Okayama, Japan, reports that low-dispersion spectra (range 440–720 nm; resolution ∼ 1 nm) of this object, obtained on Apr. 18.80 UT with the Fujii Bisei Observatory 0.28-m telescope, show a strong Hα emission line with a width of 2800 km/s (FWHM). Other emission lines include the Fe II series (37, 38, 42, 49, 55, 74), Na I (589 nm), and He I (706.5 nm), characteristic of an Fe II-class nova. Additional photometry provided by T. Kato, Kyoto: Apr. 4.769, [11.8 (A. Takao, Kita-kyushu, Japan, unfiltered CCD); 12.794, [14.0 (K. Haseda, Aichi, Japan, T-Max 400 film; revision to IAUC 7879); 17.828, 9.6 (Takao); 17.769, V = 10.16 (K. Nakajima, Kumano, Mie, Japan, CCD); 17.788, R_{c} = 9.31 (Nakajima).

COMETS C/2002 G4, C/2002 G5, C/2002 H1 (SOHO) Further to IAUC 7875, D. Hammer has reported his measurements of

three additional Kreutz sungrazers. C/2002 G4, C/2002 G5, and C/2002 H1 were found by X. Leprette, R. Kracht, and K. Černis, respectively, on SOHO website C2 images. C/2002 H1 was also visible in the C3 coronagraph. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 G4 Apr. 4.479  1^{h}01.^{m}3  + 5^{o}13′  2002-H15
C/2002 G5  12.826  1 32.1  + 8 01  2002-H15
C/2002 H1  17.113  1 58.0  + 7 58  2002-H26
2002 April 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green